According to the Guinness Book of World Records, the longest home run ever measured was hit by Roy "Dizzy"Carlyle in a minor league game. The ball traveled 196.4 meters before landing on the ground outside the ballpark. Assuming the ball's initial velocity was in a direction 45° above the horizontal and ignoring air resistance, what did the initial speed of the ball,in meters per second, need to be to produce such a home run if the ball was hit at a point 0.7 m above the ground level? How far would the ball be above a fence 3.0 meters high if the fence was 98 meters from the home plate? Write your answer in meters.


1
Expert's answer
2022-03-06T18:06:23-0500

"v=\sqrt{\frac{gl}{\sin2\alpha}}=43.8~\frac ms."